OpenStudy (anonymous):

carol bought a new winter coat at 35% off and paid $156. what was the original price of the coat?

OpenStudy (anonymous):

$156=65% of Price 156=65/100 * Price Price = 15000/65 = 240 \[P=240\]

OpenStudy (anonymous):

Thanks word problems are just so confusing to me.